# Fast Tensor Completion via Approximate Richardson Iteration

### v1

@@ Abstract
We study tensor completion (TC) through the lens of low-rank tensor decomposition (TD). Many TD algorithms use fast alternating minimization methods, which solve highly structured linear regression problems at each step (e.g., for CP, Tucker, and tensor-train decompositions). However, such algebraic structure is lost in TC regression problems, making direct extensions unclear. To address this, we propose a lifting approach that approximately solves TC regression problems using structured TD regression algorithms as blackbox subroutines, enabling sublinear-time methods. We theoretically analyze the convergence rate of our approximate Richardson iteration-based algorithm, and we demonstrate on real-world tensors that its running time can be 100x faster than direct methods for CP completion.
@@ 1 Introduction
Tensor completion (TC) is the higher-order generalization of matrix completion. It has many applications across data mining, machine learning, signal processing, and statistics (see Song et al. (2019) for a detailed survey of applications). In this problem, we are given a partially observed tensor as input (i.e., only a subset of its entries is known), and the goal is to impute the missing values. Under low-rank and other statistical assumptions, we can hope to recover the missing values by minimizing a loss function over only the observed entries.
@@
where $\bm{\theta}$ are the learnable parameters, $\widehat{\bm{\mathscr{X}}}(\bm{\theta})\in\mathbb{R}^{I_{1}\times\cdots\times I_{N}}$ is the reconstructed tensor, $\mathcal{L}_{\Omega}$ is the loss defining the error between $\widehat{\bm{\mathscr{X}}}(\bm{\theta})$ and $\bm{\mathscr{X}}$ on entries in $\Omega$, and $\mathcal{R}(\bm{\theta})$ is the regularization term. For brevity, we write $\widehat{\bm{\mathscr{X}}}$ without explicit dependence on $\bm{\theta}$. Rank constraints can be incorporated into (1) by including appropriate penalty terms in $\mathcal{R}(\bm{\theta})$.
@@
To introduce our approach, we use a running example where $\widehat{\bm{\mathscr{X}}}=\bm{\mathscr{G}}\times_{1}\mathbf{A}^{(1)}\times_{2}\mathbf{A}^{(2)}\cdots\times_{N}\mathbf{A}^{(N)}$ is a low-rank Tucker decomposition with core tensor $\bm{\mathscr{G}}\in\mathbb{R}^{R_{1}\times\dots\times R_{N}}$ and factor matrices $\mathbf{A}^{(n)}\in\mathbb{R}^{I_{n}\times R_{n}}$, and ×$\times_{i}$ is the mode-$i$ product.
In general, TC is a non-convex and NP-hard problem (Hillar and Lim, 2013). The special case where $\Omega=[I_{1}]\times\cdots\times[I_{N}]$ is the widely studied tensor decomposition (TD) problem, for which alternating least squares (ALS) methods are regularly used to compute $\bm{\mathscr{G}},\mathbf{A}^{(1)},\ldots,\mathbf{A}^{(N)}$.
@@
where $\otimes$ is the Kronecker product and $\operatorname*{vec}(\bm{\mathscr{G}}^{\prime})\in\mathbb{R}^{R}$ is the flattened version of $\bm{\mathscr{G}}^{\prime}$ with $R:=\prod_{n=1}^{N}R_{n}$.
Diao et al. (2019) and Fahrbach et al. (2022) recently exploited the Kronecker product structure in (3) to design algorithms with running times that are sublinear in the size of the full tensor $I:=\prod_{n=1}^{N}I_{n}$ by using leverage score sampling to approximately solve (3).
@@
A natural idea to overcome the lack of structure in the $\Omega$-masked updates is to lift (4) to a higher-dimensional problem by introducing variables $\mathbf{b}_{\overline{\Omega}}$, where $\overline{\Omega}$ is the complement of $\Omega$, i.e., letting the unobserved entries in $\bm{\mathscr{X}}$ be free variables. This is a convex problem with much of the same structure as the design matrix in the full TD problem, and gives the same solution as the TC update in (4):
@@
To our knowledge, such ideas date back to Healy and Westmacott (1956) in the experimental design and causal inference literature. However, it is not clear a priori that lifting is helpful for computational efficiency—it restores the structure of the design matrix, but the new problem (5) is larger and higher dimensional. Our work proposes to solve (5) with the following two-step procedure, called mini-ALS. Given $\mathbf{x}^{(k)}$, it repeats the following:
@@
    Set $\mathbf{b}^{(k)}_{\overline{\Omega}}\leftarrow\mathbf{A}_{\overline{\Omega}}\mathbf{x}^{(k)}$ and $\mathbf{b}^{(k)}_{\Omega}\leftarrow\mathbf{b}_{\Omega}$.
@@
    Set $\mathbf{x}^{(k+1)}\leftarrow\operatorname*{arg\,min}_{\mathbf{x}}\lVert\mathbf{A}\mathbf{x}-\mathbf{b}^{(k)}\rVert_{2}$.
Mini-ALS iterations can still be expensive since ${\mathbf{A}\in\mathbb{R}^{I\times R}}$ is a tall-and-skinny matrix (i.e., $I\gg R$). However, step two is a structured ALS for TD, so we propose solving it approximately with row sampling, allowing us to tap into a rich line of work on leverage score sampling for CP decomposition (Cheng et al., 2016; Larsen and Kolda, 2022; Bharadwaj et al., 2023), Tucker decomposition (Diao et al., 2019; Fahrbach et al., 2022), and tensor-train decomposition (Bharadwaj et al., 2024). Returning to the running Tucker completion example, updating the core tensor in step two of mini-ALS via (3) only requires sampling $\tilde{O}(R)$ rows of $\mathbf{A}$, which leads to a substantially smaller problem. Further, we can compute $\mathbf{b}^{(k)}$ lazily, i.e., only the entries corresponding to sampled rows. We call our lifted iterative method approximate-mini-ALS.
@@ 1.1 Our Contributions
  • In Section 3, we propose using the mini-ALS algorithm for each step of ALS for TC, and show that it simulates the preconditioned Richardson iteration (Lemma 3.5). Our main theoretical contribution is proving that the second step of a mini-ALS iteration can be performed approximately. We quantify how small this approximation error must be for approximate-mini-ALS to converge at the same rate as the Richardson iteration (Theorem 3.7). This lets us to extend a recent line of work on leverage score sampling-based TD ALS algorithms to the TC setting.
@@
  • In Section 5, we compare the empirical performance of our lifted algorithm to direct methods for a low-rank CP completion task on synthetic and real-world tensors. We observe that mini-ALS can be orders of magnitude faster than direct ALS methods, while achieving comparable reconstruction errors. Finally, we propose an accelerated version with adaptive step sizes that extrapolates the trajectory of $\mathbf{x}^{(k)}$ and converges in fewer total iterations.
@@ 1.2 Related Work
Tomasi and Bro (2005) proposed an ALS algorithm for CP completion that repeats the following two-step process: (1) fill in the missing values using the current CP decomposition, and (2) update one factor matrix. Their algorithm is equivalent to running one iteration of mini-ALS in each step of ALS. As Tomasi and Bro (2005) discuss, this can lead to slower convergence and an increased likelihood of converging to suboptimal local minima because of errors introduced by the imputed missing values.
In contrast, approximate-mini-ALS runs until convergence in each step of ALS. By doing this, we establish a connection to the Richardson iteration and build on its convergence guarantees. Further, Tomasi and Bro (2005) explicitly fill in all missing values using the current decomposition, whereas we only impute missing values required by row sampling, allowing us to achieve sublinear-time updates in the size of the tensor. In general, iteratively fitting to imputed missing values falls under the umbrella of expectation-maximization (EM) algorithms (Little and Rubin, 2019, Chapter 8).
@@
Similar to minimizing the nuclear norm for matrix completion (Fazel, 2002; Candes and Recht, 2012), a line of research in noisy TC suggests minimizing a convex relaxation of rank and identifies statistical assumptions under which the problem is recoverable (Barak and Moitra, 2016). Two standard assumptions are incoherence and the missing-at-random assumption. In Section 3, we discuss how these assumptions provide a bound on the number of steps required for mini-ALS. Alternating minimization approaches have also been applied in this line of noisy TC research (Jain and Oh, 2014; Liu and Moitra, 2020). It is also consistently observed that methods based on TD and tensor unfoldings are more practical and computationally efficient (Acar et al., 2011; Montanari and Sun, 2018; Filipović and Jukić, 2015; Shah and Yu, 2019, 2023).
@@ 2 Preliminaries
The order of a tensor is the number of its dimensions $N$. We denote scalars by normal lowercase letters $x\in\mathbb{R}$, vectors by boldface lowercase letters $\mathbf{x}\in\mathbb{R}^{n}$, matrices by boldface uppercase letters $\mathbf{X}\in\mathbb{R}^{m\times n}$, and higher-order tensors by boldface script letters $\bm{\mathscr{X}}\in\mathbb{R}^{I_{1}\times\dots\times I_{N}}$. We use normal uppercase letters for the size of an index set, e.g., $[N]=\{1,2,\dots,N\}$. We define $I_{\neq k}:=I/I_{k}$ for $k\in[N]$ and similarly $R_{\neq k:=R/R_{k}}$. We denote the $i$-th entry of $\mathbf{x}$ by $x_{i}$, the $(i,j)$-th entry of $\mathbf{X}$ by $x_{ij}$, and the $(i,j,k)$-th entry of a third-order tensor $\bm{\mathscr{X}}$ by $x_{ijk}$.
@@ 2.1 Tensor Decompositions
The tensor decompositions of $\bm{\mathscr{X}}\in\mathbb{R}^{I_{1}\times\dots\times I_{N}}$ below can be viewed as higher-order analogs of low-rank matrix factorization. See Kolda and Bader (2009) for a comprehensive survey on this topic.
@@
A rank-$\mathbf{r}$ Tucker decomposition represents $\bm{\mathscr{X}}$ as a core tensor $\bm{\mathscr{G}}\in\mathbb{R}^{R_{1}\times\dots\times R_{N}}$ and $N$ factor matrices $\mathbf{A}^{(n)}\in\mathbb{R}^{I_{n}\times R_{n}}$, for $n\in[N]$, where $\mathbf{r}=(R_{1},\dots,R_{N})$ is the multilinear rank (Ghadiri et al., 2023a). The reconstructed tensor $\widehat{\bm{\mathscr{X}}}=\bm{\mathscr{G}}\times_{1}\mathbf{A}^{(1)}\times_{2}\dots\times_{N}\mathbf{A}^{(N)}$ is defined elementwise as:
@@
A rank-$\mathbf{r}$ tensor train (TT) decomposition (Oseledets, 2011) represents $\bm{\mathscr{X}}$ using $N$ third-order TT-cores $\bm{\mathscr{A}}^{(n)}\in\mathbb{R}^{R_{n-1}\times I_{n}\times R_{n}}$, for $n\in[N]$, with the convention $R_{0}=R_{N}=1$. The reconstructed tensor $\widehat{\bm{\mathscr{X}}}$ is defined elementwise as:
@@ 2.2 ALS Formulations
Alternating least squares (ALS) methods are the gold standard for low-rank tensor decomposition, e.g., they are the first techniques mentioned in the MATLAB Tensor Toolbox (Bader and Kolda, 2023). ALS cyclically minimizes the original least-squares problem (1) with respect to one block of factor variables while keeping all others fixed. Repeating this process converges to a nontrivial local optimum and reduces the original nonconvex problem to a series of (convex) linear regression problems in each step (see Section A.1).
Updating a block of variables in an ALS step for a TD problem is typically a highly structured regression problem that can be solved very fast with problem-specific algorithms. We describe the induced structure of each ALS update for the tensor decomposition types above.
@@
The core-tensor ALS update solves: with all factors $\mathbf{A}^{(n)}$ fixed,
@@
where $\mathbf{G}_{(n)}$, $\mathbf{X}_{(n)}$ are the mode-$n$ unfoldings of $\bm{\mathscr{G}}$ and $\bm{\mathscr{X}}$. This is equivalent to solving $I_{n}$ independent linear regression problems, where the design matrix is the product of a Kronecker product and another matrix. It can also be viewed as solving $I_{n}$ instances of structured but constrained linear regression (Fahrbach et al., 2022).
@@
Given a TT decomposition $\{\bm{\mathscr{A}}^{(n)}\}_{n=1}^{N}$ and $n\in[N]$, the left chain $\mathbf{A}_{<n}\in\mathbb{R}^{(I_{1}\cdots I_{n-1})\times R_{n-1}}$ and right chain $\mathbf{A}_{>n}\in\mathbb{R}^{R_{n}\times(I_{n+1}\cdots I_{N})}$ are matrices depending on the cores $\bm{\mathscr{A}}^{(n^{\prime})}$ for $n^{\prime}<n$ and $n^{\prime}>n$, respectively (see Section A.3 for details). When ALS updates $\bm{\mathscr{A}}^{(n)}$ with all other TT-cores fixed, it solves

### v2

@@ Abstract
We study tensor completion (TC) through the lens of low-rank tensor decomposition (TD). Many TD algorithms use fast alternating minimization methods to solve highly structured linear regression problems at each step (e.g., for CP, Tucker, and tensor-train decompositions). However, such algebraic structure is often lost in TC regression problems, making direct extensions unclear. This work proposes a novel lifting method for approximately solving TC regression problems using structured TD regression algorithms as blackbox subroutines, enabling sublinear-time methods. We analyze the convergence rate of our approximate Richardson iteration-based algorithm, and our empirical study shows that it can be 100x faster than direct methods for CP completion on real-world tensors.
Tensor completion
@@ 1 Introduction
Tensor completion (TC) is the higher-order generalization of matrix completion. It has many applications across data mining, machine learning, signal processing, and statistics (see Song et al. (2019) for a detailed survey of applications). In the TC problem, we are given a partially observed tensor as input (i.e., only a subset of its entries is known), and the goal is to impute the missing values. Under low-rank and other statistical assumptions, we can recover the missing values by minimizing a loss function over only the observed entries.
@@
where $\bm{\theta}$ are the learnable parameters, $\widehat{\bm{\mathscr{X}}}(\bm{\theta})\in\mathbb{R}^{I_{1}\times\cdots\times I_{N}}$ is the reconstructed tensor, $\mathcal{L}_{\Omega}$ is the loss function defining the error between $\widehat{\bm{\mathscr{X}}}(\bm{\theta})$ and $\bm{\mathscr{X}}$ on entries in $\Omega$, and $\mathcal{R}(\bm{\theta})$ is the regularization term. For brevity, we write $\widehat{\bm{\mathscr{X}}}$ without explicit dependence on $\bm{\theta}$. Rank constraints can be incorporated into (1) by including appropriate penalty terms in $\mathcal{R}(\bm{\theta})$, e.g., by using $\ell_{1}$ or $\ell_{2}$ regularization terms to reduce the effective dimension (Fahrbach et al., 2022).
@@
To introduce our approach, we use a running example where $\widehat{\bm{\mathscr{X}}}=\bm{\mathscr{G}}\times_{1}\mathbf{A}^{(1)}\times_{2}\mathbf{A}^{(2)}\cdots\times_{N}\mathbf{A}^{(N)}$ is a low-rank Tucker decomposition with core tensor $\bm{\mathscr{G}}\in\mathbb{R}^{R_{1}\times\dots\times R_{N}}$ and factor matrices $\mathbf{A}^{(n)}\in\mathbb{R}^{I_{n}\times R_{n}}$. We do not focus on regularization in this paper, but our approach extends to regularized problems, provided there exists an algorithm that accelerates the corresponding structured and regularized objectives. For example, Fahrbach et al. (2022) present subquadratic‑time algorithms for Kronecker ridge regression (i.e., $\ell_{2}$ regularization).
In general, TC is a nonconvex and NP-hard problem (Hillar & Lim, 2013). The special case where $\Omega=[I_{1}]\times\cdots\times[I_{N}]$ is the widely studied tensor decomposition (TD) problem, for which alternating least squares (ALS) methods are often used to compute $\bm{\mathscr{G}},\mathbf{A}^{(1)},\ldots,\mathbf{A}^{(N)}$.
@@
where $\otimes$ is the Kronecker product and $\operatorname*{vec}(\bm{\mathscr{G}}^{\prime})\in\mathbb{R}^{R}$ is the flattened version of tensor $\bm{\mathscr{G}}^{\prime}$, with $R:=\prod_{n=1}^{N}R_{n}$.
Diao et al. (2019) and Fahrbach et al. (2022) recently exploited the Kronecker product structure in (3) to give algorithms with running times that are sublinear in the size of the full tensor $I:=\prod_{n=1}^{N}I_{n}$. These approaches use leverage score sampling to approximately solve (3).
@@
A natural idea to overcome the lack of structure in the $\Omega$-masked updates is to lift (4) to a higher-dimensional problem by introducing variables $\mathbf{b}_{\overline{\Omega}}$, where $\overline{\Omega}$ is the complement of $\Omega$, i.e., letting the unobserved entries in $\bm{\mathscr{X}}$ be free variables. This is a convex problem with much of the same structure as the design matrix in the full TD problem. Further, it gives the same solution as the TC update in (4):
@@
To our knowledge, such ideas date back to Healy & Westmacott (1956) in the experimental design and causal inference literature. However, it is not clear a priori that lifting is helpful for computational efficiency—it restores the structure of the design matrix, but the new problem (5) is larger and higher dimensional.
This work proposes solving (5) with a two-step procedure called mini-ALS. Given a vector $\mathbf{x}^{(k)}$ corresponding to an iterate of a block of variables in the low-rank decomposition, it repeats the following:
@@
    Set $\mathbf{b}^{(k)}_{\overline{\Omega}}\leftarrow\mathbf{A}_{\overline{\Omega}}\mathbf{x}^{(k)}$ and $\mathbf{b}^{(k)}_{\Omega}\leftarrow\mathbf{b}_{\Omega}$ // lift
@@
    Set $\mathbf{x}^{(k+1)}\leftarrow\operatorname*{arg\,min}_{\mathbf{x}}\lVert\mathbf{A}\mathbf{x}-\mathbf{b}^{(k)}\rVert_{2}$ // solve
Mini-ALS iterations can still be expensive since $\mathbf{A}\in\mathbb{R}^{I\times R}$} is a tall-and-skinny matrix (i.e., $I\gg R$). However, step two is a structured ALS for TD, so we propose solving it approximately with row sampling, allowing us to tap into a rich line of work on leverage score sampling for CP decomposition (Cheng et al., 2016; Larsen & Kolda, 2022; Bharadwaj et al., 2023), Tucker decomposition (Diao et al., 2019; Fahrbach et al., 2022), and tensor-train decomposition (Bharadwaj et al., 2024).
Returning to the running Tucker completion example, updating the core tensor in step two of mini-ALS via (3) only requires sampling $\tilde{O}(R)$ rows of $\mathbf{A}$, which is a substantially smaller problem. Further, we can compute $\mathbf{b}^{(k)}$ lazily, i.e., only the entries corresponding to sampled rows. We call our lifted iterative method approximate-mini-ALS.
@@ 1.1 Our Contributions
  • In Section 3, we propose using the mini-ALS algorithm for each step of ALS for TC. We show that it simulates the preconditioned Richardson iteration (Lemma 3.5). Our main theoretical contribution is proving that the second step of a mini-ALS iteration can be performed approximately. We quantify how small the approximation error must be for approximate-mini-ALS to converge at the same rate as the Richardson iteration (Theorem 3.7). This lets us to extend a recent line of work on leverage score sampling-based TD ALS algorithms to the TC setting.
@@
  • In Section 5, we show that leverage score sampling is an effective method for solving large structured regression problems via the coupled matrix problem. Then we compare the empirical performance of our lifted algorithm to direct methods for low-rank CP completion on synthetic and real-world tensors. We observe that mini-ALS can be orders of magnitude faster than direct ALS methods, while achieving comparable reconstruction errors. Finally, we propose an accelerated version with adaptive step sizes that extrapolates the trajectory of the iterates $\mathbf{x}^{(k)}$ and converges in fewer iterations.
@@ 1.2 Related Work
Tomasi & Bro (2005) proposed an ALS algorithm for CP completion that repeats the following two-step process: (1) fill in the missing values using the current CP decomposition, and (2) update one factor matrix. Their algorithm is equivalent to running one iteration of mini-ALS in each step of ALS. As Tomasi & Bro (2005) discuss, this can lead to slower convergence and an increased likelihood of converging to suboptimal local minima because of errors introduced by the imputed missing values.
In contrast, approximate-mini-ALS runs until convergence in each step of ALS. By doing this, we establish a connection to the Richardson iteration and build on its convergence guarantees. Further, Tomasi & Bro (2005) explicitly fill in all missing values using the current decomposition, whereas we only impute missing values required by row sampling, allowing us to achieve sublinear-time updates in the size of the tensor. In general, iteratively fitting to imputed missing values falls under the umbrella of expectation-maximization (EM) algorithms (Little & Rubin, 2019, Chapter 8).
@@
Similar to minimizing the nuclear norm for matrix completion (Fazel, 2002; Candes & Recht, 2012), a line of research in noisy TC proposes minimizing a convex relaxation of rank and identifies statistical assumptions under which the problem is recoverable (Barak & Moitra, 2016). Two standard assumptions are incoherence and the missing-at-random assumption. In Section 3, we discuss how these assumptions provide a bound on the number of steps required for mini-ALS. Alternating minimization approaches have also been applied in this line of noisy TC research (Jain & Oh, 2014; Liu & Moitra, 2020). It is also consistently observed that methods based on TD and tensor unfoldings are more practical and computationally efficient (Acar et al., 2011; Montanari & Sun, 2018; Filipović & Jukić, 2015; Shah & Yu, 2019, 2023).
@@ 2 Preliminaries
The order of a tensor is the number of its dimensions $N$. We denote scalars by normal lowercase letters $x\in\mathbb{R}$, vectors by boldface lowercase letters $\mathbf{x}\in\mathbb{R}^{n}$, matrices by boldface uppercase letters $\mathbf{X}\in\mathbb{R}^{m\times n}$, and higher-order tensors by boldface script letters $\bm{\mathscr{X}}\in\mathbb{R}^{I_{1}\times\dots\times I_{N}}$. We use normal uppercase letters for the size of an index set, e.g., $N]=\{1,2,\dots,N\}$. We define $I_{\neq k}:=I/I_{k}$ for $k\in[N]$ and similarly $R_{\neq k:=R/R_{k}}$. We denote the $i$-th entry of $\mathbf{x}$ by $x_{i}$, the $(i,j)$-th entry of $\mathbf{X}$ by $x_{ij}$, and the $(i,j,k)$-th entry of a third-order tensor $\bm{\mathscr{X}}$ by $x_{ijk}$.
@@ 2.1 Tensor Decompositions
The tensor decompositions of $\bm{\mathscr{X}}\in\mathbb{R}^{I_{1}\times\dots\times I_{N}}$ below can be seen as higher-order analogs of low-rank matrix factorization. We direct the reader to Kolda & Bader (2009) for a comprehensive survey on this topic.
@@
A rank-$\mathbf{r}$ Tucker decomposition represents $\bm{\mathscr{X}}$ with a core tensor $\bm{\mathscr{G}}\in\mathbb{R}^{R_{1}\times\dots\times R_{N}}$ and $N$ factor matrices $\mathbf{A}^{(n)}\in\mathbb{R}^{I_{n}\times R_{n}}$, for $n\in[N]$, where $\mathbf{r}=(R_{1},\dots,R_{N})$ is the multilinear rank (Ghadiri et al., 2023a). The reconstructed tensor $\widehat{\bm{\mathscr{X}}}=\bm{\mathscr{G}}\times_{1}\mathbf{A}^{(1)}\times_{2}\dots\times_{N}\mathbf{A}^{(N)}$ is defined elementwise as:
@@
A rank-$\mathbf{r}$ tensor train (TT) decomposition (Oseledets, 2011) represents $\bm{\mathscr{X}}$ with $N$ third-order TT-cores $\bm{\mathscr{A}}^{(n)}\in\mathbb{R}^{R_{n-1}\times I_{n}\times R_{n}}$, for $n\in[N]$, using the convention $R_{0}=R_{N}=1$. The reconstructed tensor $\widehat{\bm{\mathscr{X}}}$ is defined elementwise as:
@@ 2.2 ALS Formulations
Alternating least squares (ALS) methods are the gold standard for low-rank tensor decomposition, e.g., they are the first techniques mentioned in the MATLAB Tensor Toolbox (Bader & Kolda, 2023). ALS cyclically minimizes the original least-squares problem (1) with respect to one block of factor variables while keeping all others fixed. Repeating this process converges to a nontrivial local optimum and reduces the original nonconvex problem to a series of (convex) linear regression problems in each step (see Section A.1).
Updating a block of variables in an ALS step of a TD problem is often a highly structured regression problem that can be solved very fast with problem-specific algorithms. We describe the induced structure of each ALS update for the tensor decomposition types above.
@@
The core-tensor ALS update solves the following: for $\mathbf{A}^{(n)}$ fixed,
@@
where $\mathbf{G}_{(n)}$, $\mathbf{X}_{(n)}$ are the mode-$n$ unfoldings of $\bm{\mathscr{G}}$ and $\bm{\mathscr{X}}$. This is equivalent to solving $I_{n}$ independent linear regression problems, where the design matrix is the product of a Kronecker product and another matrix. It can be viewed as solving $I_{n}$ instances of structured but constrained linear regression (Fahrbach et al., 2022).
@@
Given a TT decomposition $\{\bm{\mathscr{A}}^{(n)}\}_{n=1}^{N}$ and $n\in[N]$, the left chain $\mathbf{A}_{<n}\in\mathbb{R}^{(I_{1}\cdots I_{n-1})\times R_{n-1}}$ and right chain $\mathbf{A}_{>n}\in\mathbb{R}^{R_{n}\times(I_{n+1}\cdots I_{N})}$ are matrices that depend on the cores $\bm{\mathscr{A}}^{(n^{\prime})}$ for $n^{\prime}<n$ and $n^{\prime}>n$, respectively (see Section A.3 for details). When ALS updates $\bm{\mathscr{A}}^{(n)}$ with all other TT-cores fixed, it solves